Show that the test taking the overall significance of regression model using ANOVA table to be expressed as:Β 

Β 

π‘ΉπŸβ„π’Œβˆ’πŸ

Β 

𝑭=(πŸβˆ’π‘ΉπŸ)β„π’βˆ’π’ŒΒ 

Β 

Where, R be a level of determination and k is the number of parameters in the n sampled regression model.


1
Expert's answer
2021-05-31T15:56:35-0400

"F=\frac{MSR}{MSE}"

Where;

"MSR= \frac{SSR}{1}"

And

"MSE = \frac{SSE}{n-2}"

"F=\frac{1-R^2}{n-k}"